### v2

# Examining dark neutron decay using f-mode gravitational waves

Abstract
In this study, the impact of neutron decay into dark matter and various dark matter self-interaction strengths on neutron star properties have been explored. Using the quark-meson coupling (QMC) model for nucleon-only equations of state (EoSs), the effects of different matter compositions have been compared, including strange matter and self-interacting dark matter. The results demonstrate that increasing DM-DM self-repulsion stiffens the EoS, influencing the mass-radius relationship and stability of neutron stars.
Furthermore, fundamental mode (f-mode) oscillations have been analyzed, which serve as a diagnostic tool for probing neutron star interiors. The f-mode frequencies follow universal relations, reinforcing their applicability for constraining dense matter properties. It has been shown that neutron stars composed of nucleons-only and self-interacting dark matter exhibit a universal behavior in damping time and angular frequency, whereas strange matter and non-self-interacting dark matter deviate from this trend.
These findings highlight the significance of gravitational wave detections in investigating neutron star structure, particularly with next-generation observatories such as the Einstein Telescope and Cosmic Explorer. In particular, the f-mode frequencies generated by such stars fall within the detection range of modern telescopes, making it possible to observe these signals directly. This opens a promising pathway for probing dark matter indirectly—using neutron stars as natural laboratories for fundamental physics.
1 Introduction
The hypothesis of self-interacting dark matter (SIDM) has gained prominence due to its potential to address discrepancies within the standard collisionless cold dark matter (CCDM) paradigm, alongside its theoretical plausibility within extensions of the Standard Model. Specifically, the CCDM model faces challenges such as the observed flat density profiles in dwarf galaxy cores, contrasting with the cuspy profiles predicted by simulations [1, 2], the overabundance of simulated satellite galaxies compared to observed counts [1, 3, 4], and the ”too big to fail” problem, wherein simulations predict massive, star-forming dwarf galaxies that remain undetected [5, 6].
While baryonic interactions and statistical fluctuations have been proposed as potential resolutions to galactic anomalies, SIDM offers a compelling alternative capable of addressing these issues, and the neutron decay hypothesis provides a specific mechanism for dark matter production. Enhanced self-scattering rates in high-density regions, resulting from SIDM, could effectively flatten dwarf galaxy cores. The required self-interaction cross-sections, within the range of $^{2}/g<\sigma_{\chi\chi}/m_{\chi}<10 cm$$^{2}/g$, where $\sigma_{\chi\chi}$ is the self-interaction cross-section and $m_{\chi}$ the dark matter particle mass are, have been extensively studied [7, 8, 9, 10, 11].
Additionally, the neutron lifetime anomaly, where beam experiments yield longer lifetimes than bottle experiments [12, 13, 14, 15], suggests a potential decay channel into undetected dark sector particles [16, 17, 18, 19]. This has motivated the exploration of neutron decay into dark matter as a potential resolution. The neutron decay scenario, proposed by Fornal and Grinstein [18], posits that neutrons decay into a dark fermion ($\chi$), potentially explaining the observed neutron lifetime discrepancy. This hypothesis implies that neutron stars, composed primarily of neutrons, could serve as natural laboratories for studying dark matter interactions.
The fundamental mode (f-mode) is a type of non-radial oscillation in compact stars, primarily governed by the star’s bulk properties—such as mass, radius, and internal composition. These oscillations generate gravitational waves, and their frequency and damping time are sensitive to the equation of state (EoS) of the stellar interior. As such, the f-mode serves as a powerful diagnostic tool for probing the internal structure and composition of neutron stars and, more generally, for constraining the nature of dense matter under extreme conditions.
There are several promising avenues for measuring f-mode [20, 21, 22, 23] characteristics in compact stars. For instance, third-generation gravitational wave detectors like the Einstein Telescope [24] and Cosmic Explorer [25] are expected to achieve the sensitivity required for detecting f-mode signals from astrophysical sources [26].
Furthermore, the damping time and frequency of the f-mode are also linked to the star’s properties, such as mass, compactness, and moment of inertia, quantities that could be inferred from spin-orbit coupling effects in binary pulsar systems [27]. These interdependencies allow the f-mode to act as a bridge between observable gravitational wave signatures and the otherwise inaccessible interior physics of compact stars.
In this study, the neutron decay into dark matter particle model is tested using well-established universal relations for neutron stars (NSs) involving f-mode properties—relations that have been foundational in the field due to their near-independence from the specific choice of nuclear equation of state (EoS), and which have recently been revisited [28, 29]. This investigation is further extended to isolated compact stars composed of dark matter (DM), modeled as static, spherically symmetric objects with EoSs for both fermionic and bosonic dark matter particles.
This paper has been divided into different sections. The layout of this study is as follows: Section 2 provides theoretical possibility of neutron decay into dark matter particles, examining potential mechanisms and astrophysical constraints. Section 3 provides a comprehensive overview of neutron stars, detailing their formation, key properties, and observed phenomena. Section 4 focuses on the crucial concept of the Equation of State (EoS) governing the ultradense matter within neutron stars, exploring theoretical models and their implications. Building upon this, Section 5 discusses the fundamental structural equations that describe the equilibrium of neutron stars under their immense gravity. Section 6 further elaborates on the Equation of State, potentially exploring specific aspects or advanced considerations. Section 7 analyzes the intriguing phenomenon of universalities observed in neutron star properties, highlighting equation-of-state independent relationships. This section will be further subdivided, with Section 7.1 specifically focusing on the damping time of neutron star oscillations and its physical mechanisms. Section 7.2 will then examine the mass-scaled angular frequency of rotating neutron stars and its significance. Section 8 provides a comparative analysis of f-mode frequencies across various neutron star models, offering insights into their structural and dynamical characteristics. Section 9 concludes the article by summarizing the key findings and discussing their broader relevance to neutron star physics and gravitational wave astronomy.
2 Neutron decay into dark matter
There remains an ongoing discrepancy between neutron lifetime measurements obtained using beam and bottle methods. Beam experiments, which detect proton decay, consistently yield lifetimes approximately 1% longer than those measured in bottle experiments, where the decay products are not directly observed. Advances in ultracold neutron studies have refined bottle measurements, yet beam experiments continue to produce lifetimes that remain about eight seconds longer [30, 19, 31, 32, 33, 34, 35, 36, 37, 38, 37].
To address this anomaly, Fornal and Grinstein [16, 17, 18] proposed a dark decay channel wherein neutrons decay into dark sector particles,
$n\xrightarrow{}\chi+\phi$
including a dark fermion ($\chi$), alongside the standard mode.
This hypothesis posits that weakly interacting dark matter particles remain undetected in beam experiments but contribute to the total decay rate in bottle experiments. This alternative to neutron-mirror neutron oscillation [16, 17] necessitates a dark fermion with a mass (mχ) nearly degenerate with the neutron within the range $937.9\,\text{MeV}<m_{\chi}<938.7\,\text{MeV}$, more details of the model can be found in Ref. [16, 17]. The accompanying boson cannot be a photon as shown in Ref.  [39, 40].
Furthermore, this dark decay hypothesis has been suggested as a potential resolution to the reactor anti-neutrino anomaly [40]. Notably, it imposes constraints on neutron star properties. If these dark fermions are non-interacting, the maximum neutron star mass is predicted to be below 0.7 M [41, 42, 43, 44, 45], contradicting observed neutron star masses exceeding 2 M [46, 47, 48, 49]. Therefore, a repulsive self-interaction among dark fermions is required to reconcile this discrepancy.
This study investigates the implications of this dark decay channel on neutron star properties. This study takes forward the previous findings and constrains [50, 17, 51, 42, 16] the properties of the dark matter using the gravitational wave, which can be very crucial in determining the interior of the neutron star.
3 Neutron Stars
Neutron stars are extremely compact objects that exist in the universe and exhibit an extreme range of energy densities. In fact, black holes are the only objects that are more compact than neutron stars. The energy densities at the outer crust are comparatively lower, and this part plays a crucial role in determining the radius of the neutron stars At the core, neutron stars contain energy densities several times the energy density of normal nuclear matter. The composition of matter at such high energy densities has been a puzzle for physicists. There have been speculations proposed by different physicists suggesting such high energy densities at the core of neutron stars may contain nuclear matter, hyperons, or even strange matter [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. But it has been widely accepted that mostly 90% of it is made of neutrons. This neutron-rich environment renders neutron stars ideal astrophysical laboratories for examining the proposed neutron decay into dark matter (DM) scenarios.
The hypothesis that neutrons may decay into DM suggests that a substantial number of dark matter particles may accumulate within neutron stars, assuming dark matter is annihilating, potentially leading to observable alterations in their macroscopic properties [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 44, 97, 98, 99, 100, 101, 102, 103, 104]. To model the nuclear matter inside neutron stars with precision, the equation of state (EoS) based on the quark-meson coupling (QMC) model has been used [105, 106, 107, 108]. This model is rooted in a relativistic mean field. This framework, grounded in relativistic mean-field theory, effectively represents nucleon interactions through meson exchange. In this study, is limited to neutron decay into dark matter, hadronic composition to nucleons and strange quark matter only.
4 Equation of State
To determine the equation of state, the QMC model has been adapted, which has shown promising results in [85, 50, 109, 110]. The quark-meson coupling (QMC) model was first introduced by Guichon and later expanded upon by Guichon, Thomas, and their collaborators. In this framework, nucleons are modeled as clusters of three quarks confined within an MIT bag, emphasizing their internal structure—unlike other models that treat nucleons as point-like particles [111]. The interactions between baryons arise from meson exchange, with the mesons coupling dynamically to the confined quarks. A key feature of this model is the strong scalar mean field, which induces significant modifications in the properties of bound baryons. As proposed by Fornal and Grinstein [18, 17], if neutron stars decay into dark matter. Dark matter must be taken into account when formulating an equation of state. At ($\beta$)-equilibrium, the chemical equilibrium equations governing the composition of matter are [42, 112, 76]:
$\mu_{n}=\mu_{\chi},\quad\mu_{n}=\mu_{p}+\mu_{e},\quad\mu_{\mu}=\mu_{e}$ (4.1)
The energy density of the system, including the dark matter, is given by:
$\epsilon_{H}=\frac{1}{2}m^{2}_{\sigma}\sigma^{2}+\frac{1}{2}m^{2}_{\omega}\omega^{2}+\frac{1}{2}m^{2}_{\rho}\rho^{2}+\frac{1}{\pi^{2}}\int^{k_{F}^{n}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{p}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk$ (4.2)
$+\frac{1}{\pi^{2}}\int^{k_{F}^{e}}_{0}k^{2}\sqrt{k^{2}+m_{e}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\mu}}_{0}k^{2}\sqrt{k^{2}+m_{\mu}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\chi}}_{0}k^{2}\sqrt{k^{2}+m_{\chi}^{2}}dk\,.$
where $m_{i}$ is the mass of corresponding particle (subscripted). The nucleon effective mass, $M_{N}^{*}(\sigma)$, given by [108, 109]
$M_{N}^{*}(\sigma)=M_{N}-g_{\sigma}\sigma+\frac{d\tilde{(}g_{\sigma}\sigma)^{2}}{2}\,.$ (4.3)
Here $g_{\sigma}$ is the $\sigma$-nucleon coupling constant in free space calculated in the model. The pressure ($P$) is calculated using
$P=\int\mu_{f}n_{f}-\epsilon.$ (4.4)
The nucleons-only EoS is taken from [108] and strange matter EoS is drive based on MIT bag model and the EoS used is given by
$P=\frac{1}{3}(\epsilon-4B)$ (4.5)
where $\epsilon$ the energy density of the strange matter and $B$ is the bag constant, taken to be 1014gm/fm3 [113].
5 Structural equations
Neutron stars are dense astrophysical objects shaped by intense nuclear and gravitational forces. The potential presence of dark matter within these stars may influence their structural characteristics, resulting in changes to their equation of state (EoS). This study examines methods for integrating dark neutron components into the conventional neutron star EoS and solving the Tolman-Oppenheimer-Volkoff (TOV) [114, 115] equations to assess their impact
TOV equation
$\frac{dP}{dr}=-\frac{(\epsilon+P)\frac{d\Phi}{dr}}{r}$ (5.1)
$\frac{d\Phi}{dr}=\frac{M+4\pi r^{3}P}{r(r-2M)}$ (5.2)
$\frac{dM}{dr}=4\pi r^{2}\epsilon$ (5.3)
where $P$ is the pressure, $\epsilon$ is the energy ensity, $\Phi$ is the gravitational potential, $M$ is the mass, and $r$ is the radial coordinate
Perturbation Equations for f-mode Oscillations
The f-mode oscillations are governed by two coupled differential equations [116, 117, 118]:
$\frac{dW}{dr}=\frac{d\epsilon}{dP}\left[\omega^{2}r^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W\right]-l(l+1)e^{\lambda}V$ (5.4)
$\frac{dV}{dr}=2\frac{d\Phi}{dr}V-e^{\lambda}\frac{W}{r^{2}}$ (5.5)
where $W$ represents the radial displacement function, $V$ is the perturbation of the gravitational potential, $\omega$ is the angular frequency of oscillations, and $l$ is the spherical harmonic index.
Boundary Conditions and Numerical Integration
To solve these equations, boundary conditions are applied:
  • At the center:
    $W(r)=Ar^{l+1},\quad V(r)=-Ar^{l}/l$ (5.6)
  • At the surface:
    $\omega^{2}R^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W=0$ (5.7)
The equations are integrated numerically. The eigenvalues of the system correspond to the f-mode frequencies.
6 Relationship between energy density and pressure
6.1 Effective Lagrangian for neutron–dark matter conversion
To model the possibility of neutron decay into dark matter inside neutron stars, we extend the nucleonic sector (described here using the QMC model) by including a dark fermion $\chi$ carrying baryon number. In order to reproduce the repulsive self-interaction assumed in the EoS analysis, we introduce a dark vector mediator $V_{\mu}$, analogous to the $\omega$ meson in relativistic mean-field models. The effective Lagrangian reads
$\mathcal{L}\;=\;\mathcal{L}_{\rm QMC}[n,p,\sigma,\omega_{\mu},\rho_{\mu}]+\bar{\chi}\,(i\not{\partial}-m_{\chi})\,\chi-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}-g_{\chi}\,\bar{\chi}\gamma^{\mu}\chi V_{\mu}+\varepsilon\,(\bar{n}\chi+\bar{\chi}n),$ (6.1)
where $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$. The last term represents an effective in-medium mixing between neutrons and dark fermions, which allows for neutron conversion into dark matter when $\mu_{n}>\mu_{\chi}$. Baryon number is conserved if $\chi$ carries baryon number =1=1.
At scales below $m_{V}$, integrating out the heavy vector yields a four-fermion operator,
$\mathcal{L}_{\rm eff}\supset\frac{g_{\chi}^{2}}{2m_{V}^{2}}\,(\bar{\chi}\gamma_{\mu}\chi)(\bar{\chi}\gamma^{\mu}\chi)\equiv\frac{G}{2}\,(\bar{\chi}\gamma_{\mu}\chi)^{2},$ (6.2)
where the effective coupling is defined as $G=g_{\chi}^{2}/m_{V}^{2}$. In mean-field approximation, this produces a repulsive contribution to the energy density and pressure, $\Delta\varepsilon=\tfrac{1}{2}Gn_{\chi}^{2}$, $\Delta P=\tfrac{1}{2}Gn_{\chi}^{2}$, which stiffens the equation of state as $G$ increases [41, 42, 50]. This mechanism directly explains why the dark-matter–induced EoS approaches the nucleon-only EoS for large $G$, as shown in Figure (1), which presents the equations of state (EoS) for nucleonic matter, strange matter, and scenarios involving neutron decay into dark matter under varying dark matter self-interactions. For the nucleon-only case, the Quark-Meson Coupling (QMC) model is employed. The EoS describing neutron decay into dark matter is shown for several values of the dark matter–dark matter (DM–DM) self-repulsion strength. For comparison, the strange matter EoS is also included.
The results indicate that the nucleon-only EoS is the stiffest among the considered models. As the DM–DM self-repulsion increases, the corresponding EoS for neutron decay into dark matter becomes progressively stiffer and approaches the nucleon-only case which is in agreement with the findings of [50, 41, 51]. This behavior arises because stronger DM–DM self-interaction raises the energy cost for neutron decay into dark matter, thereby reducing the overall dark matter population within the system [50]. Consequently, larger self-interaction strengths, $G$, result in fewer dark matter particles and stiffer equation of state [50].
The strange matter EoS provides a poor description of nuclear matter at low energy densities. As shown in Fig. (1), it coincides with the nucleon-only EoS up to approximately 500 MeV/fm3 [108, 109, 50]. Beyond this density, strange matter begins to manifest, and the corresponding EoS deviates from both the nucleon-only and the neutron-decay EoSs. In this regime, the strange matter EoS is comparatively softer than the other two models. The deviation of the strange matter EoS at densities above $\sim 500$ MeV/fm3 arises from the bag constant assumption in the MIT bag model, which lowers the pressure support compared to nucleonic models. As a result, the sound speed is reduced, leading to longer damping times and lower f-mode frequencies. This implies that strange stars may emit weaker gravitational waves, with potentially distinguishable signals compared to nucleonic stars of the same mass.
Figure 1: Equations of state (pressure vs. energy density) for nucleonic matter (QMC), strange matter (MIT bag model), and neutron decay into dark matter for different DM–DM self-interaction strengths $G$ (fm2). Higher $G$ corresponds to stronger repulsive self-interactions.
Fig. (2) represents the mass and radius of neutron stars based on different EoSs. Ref. [119, 120, 121] report the observation of neutron stars of a mass of exceeding 2 M [46, 47, 48, 49], indicating that any viable model must support a maximum neutron star mass above this threshold 2 M. The nucleon-only EoS produces a neutron star of maximum mass greater than 2 M, the strange matter EoS produces a neutron star of maximum mass close to 2 M, while different EoSs based on different DM-DM self-interactions produce a neutron star of less than 2 M. The non-self-interactive ($G$ = 0 fm2) DM EoS produces a neutron star of mass close to 0.7 M which is well below the acceptable value and alignes with [41, 17, 51]. When the DM-DM self-interaction $G=26\,\mathrm{fm}^{2}$ [50] the neutron decay into dark matter EoS produces a neutron star of 2 M. Therefore, acceptable self-interaction strength is $G>$ 26 fm2. The failure of the $G=0$ fm2 case to support neutron stars above $\sim 0.7M_{\odot}$ highlights that non-self-interacting dark fermions dilute pressure support too severely. Physically, the DM particles form a diffuse distribution, decreasing compactness and oscillation frequencies.
Figure 2: Mass–radius relations for neutron stars with different equations of state (EoSs). Shown are nucleon-only (QMC), strange matter (MIT bag model), and dark-matter–admixed stars with varying self-interaction strengths $G$. The horizontal band indicates the observed $2,M_{\odot}$ pulsar constraint. Increasing $G$ stiffens the dark-matter EoS, raising the maximum mass toward the nucleon-only case. The radius of light neutron stars is smaller because the EoS is not optimized in the lower energy density region (outer crust region) [109].
7 Universalities
The universalities refer to empirical relations that link the fundamental oscillation modes of compact objects, such as neutron stars and dark stars, to their macroscopic properties like mass, radius, and compactness. These universal relations are valuable because they provide a model-independent way to infer stellar characteristics from gravitational wave observations [122, 117]. Traditional neutron stars follow well-established scaling laws for f-mode frequencies and damping times, but recent studies show that dark matter admixed compact stars deviate from these relations, requiring new frameworks to describe their oscillations [123, 124]. By leveraging f-mode gravitational wave signals, astrophysicists aim to refine equations of state and probe exotic forms of matter beyond the Standard Model. The detailed analysis on tidal deformability of neutron decay mode is given in [50].
7.1 Damping time
The damping time $\tau$ of f-mode oscillations is related to the balance between the total mode energy $E$ and the rate of energy loss via gravitational-wave emission $\dot{E}{\rm{}_{GW}}$,
$\tau=\frac{2E}{\dot{E}_{GW}}.$ (7.1)
The oscillation energy for a mode with angular frequency $\omega$ is given by
$E=\frac{1}{2}\omega^{2}\int\rho(r)|\xi(r)|^{2}d^{3}r,$ (7.2)
where $\xi(r)$ is the Lagrangian displacement vector and $\rho(r)$ the local energy density. The gravitational-wave luminosity for the dominant quadrupole mode is
$\dot{E}{\rm{}_{GW}}=-\frac{32G}{5c^{5}}\omega^{6}\left|\int\rho(r)r^{2}Y{2m}(\theta,\phi)drd\Omega\right|^{2}.$ (7.3)
From these relations, the universal scaling $\tau\sim R^{4}/M^{3}$ emerges [125], which we adopt and extend in this study.
Reference [28] introduced a first-order universal relation for the damping time ($\tau$) of f-mode oscillations, derived from the quadrupole radiation formula. The relation, shown in Figure (3), is expressed as
$\tau\sim\frac{R^{4}}{M^{3}}.$ (7.4)
Here $\tau$ reflects the ratio of oscillation energy to gravitational-wave luminosity, implying that more massive stars radiate more efficiently and exhibit shorter damping times. Lioutas and Stergioulas [126] later extended this relation with higher-order corrections, confirming the monotonic decrease of $\tau$ with mass. In comparison, strange matter stars exhibit longer damping times than both nucleon-only stars and neutron stars with dark matter cores.
Figure 3: The normalised damping time of the f-mode as a function of stellar mass. Nucleon-only stars exhibit the shortest damping times, while strange-matter stars show longer-lived oscillations. Dark-matter–admixed stars interpolate between these limits depending on $G$.
7.2 Mass-scaled angular frequency
Figure 4: The relationship between mass-scaled angular frequency and compactness, where mass is given in solar masses and radius in given in kilometers, demonstrating a universal trend, is plotted for selected dark matter equations of state (EoSs). This analysis includes various DM-DM self-interaction strengths.
Figure 5: The relationship between mass-scaled angular frequency and effective compactness is plotted for selected dark matter EoSs, including various DM-DM self-interaction strengths.
Figures (4) and (5) illustrate the relation between the mass-scaled angular frequency, where $\omega_{r}=2\pi f$, and two structural parameters of neutron stars: compactness ($M/R$) and $\psi=\sqrt{M^{3}/I}$ [117]. These diagnostics are particularly valuable since they encode information about the distribution of mass and the role of internal composition in determining stellar oscillation properties. The parameter $\sqrt{M^{3}/I}$, in particular, is closely tied to the efficiency of gravitational-wave (GW) emission, as it relates the global stellar mass to the moment of inertia, thereby providing insight into how matter is distributed inside the star.
As Figure (4) shows, when $G>$ 26 fm2, the nucleon-only and self-interacting DM EoSs follow a universal scaling curve, independent of the specific interaction strength. This suggests that compactness is a reliable predictor of oscillation properties, robust against moderate variations in microphysics. Strange matter and non-self-interacting DM deviate significantly. The strange matter deviation reflects the reduced stiffness of the EoS, while the DM $G=0$ case is destabilized by the large DM population at the core, altering the star’s effective compactness. These deviations may serve as diagnostic signatures of exotic compositions.
Figure (5) further reinforces the presence of a quasi-universal relation across different microscopic models when the angular frequency is expressed in terms of $\psi=\sqrt{M^{3}/I}$. Remarkably, even when exotic compositions are considered—including strange matter and neutron-decay into DM with strong self-interaction ($G>26$ fm2)—the data points collapse onto a nearly single curve. This indicates that $\psi$ effectively captures the essential structural scaling of neutron stars, smoothing out differences in composition. Deviations are observed only in the dark matter without self-repulsion case, but the overall correlation persists. This reinforces the robustness of universal relations in gravitational-wave asteroseismology.
8 F-mode frequency
The frequency of an f-mode is highly sensitive to the EoS of the ultra-dense matter within the neutron star [116, 117, 118]. A stiffer EoS generally predicts higher f-mode frequencies. Conversely, a softer EoS predicts lower f-mode frequencies. The acceptable values of f-mode frequency for compact stars, such as neutron stars, typically range between 1 and 1–3 kHz. These frequencies are significant because they fall within the detection range of advanced gravitational wave detectors like the Einstein Telescope and Cosmic Explorer [24]. Figure (6) illustrates that the f-mode frequency of the selected models adheres to the universal relation, demonstrating consistency across different astrophysical scenarios. Notably, both frequency ranges fall within the detection capabilities of the Einstein Telescope and Cosmic Explorer, highlighting their potential for observational verification. Furthermore, for neutron decay into dark matter, variations in dark matter self-interaction strength do not disrupt the universal trend. Regardless of the DM-DM self-interaction parameter, the f-mode frequency remains well within the expected range, reinforcing the robustness of the underlying physical principles governing these oscillations. This consistency suggests that such models could be effectively probed through next-generation gravitational wave observatories.
Our results are consistent with earlier works exploring neutron-to-dark matter decay in neutron stars. Motta et al. [42] and Grinstein et al. [17] found that non-interacting dark fermions destabilize neutron stars, reducing their maximum mass below the $2M_{\odot}$ threshold, which we confirm in our analysis. Husain et al. [50] similarly emphasized the necessity of DM self-repulsion, consistent with our finding that $G>26$ fm2 restores compatibility with observational constraints.
For f-mode oscillations, our study extends the universal relations established by Andersson and Kokkotas [125] and refined by Lioutas and Stergioulas [126], showing that these relations persist even in the presence of exotic matter. However, we demonstrate that strange matter and non-self-interacting DM produce systematic deviations, suggesting that precise future detections could be used to discriminate between EoSs.
Figure 6: The f-mode frequencies are calculated for various equations of state (EoSs), taking into account different dark matter (DM) self-interaction strengths, agreeing to Ref. [127, 126].
9 Conclusion
In this work, we investigated the impact of nucleonic, strange matter, and dark matter equations of state on the f-mode oscillation properties of neutron stars, with particular attention to damping times and universal relations. Our results demonstrate that the nucleon-only EoS is consistently the stiffest, while the inclusion of dark matter generally softens the stellar structure, unless significant DM–DM self-interactions are present. Increasing the self-repulsion parameter $G$ suppresses dark matter production, thereby stiffening the EoS and bringing the neutron-decay scenario closer to the nucleonic case. In contrast, strange matter exhibits longer damping times and notable deviations at low densities, reflecting its softer pressure support relative to nucleonic matter.
Importantly, we find that neutron decay into dark matter leads to stellar configurations that remain compatible with current astrophysical constraints. Moreover, the corresponding f-mode frequencies of such stars fall squarely within the detection range of modern gravitational-wave interferometers, underscoring their observational relevance.
Despite these compositional differences, the mass-scaled angular frequency retains quasi-universal behavior when expressed as a function of compactness ($C=M/R$) and the structural parameter $\psi=\sqrt{M^{3}/I}$. This universality is preserved even in the presence of exotic matter, with only minor deviations in the case of strange matter and non-self-interacting dark matter. The persistence of such relations highlights their robustness and utility as probes of neutron star structure.
Astrophysically, these results suggest that f-mode oscillations and their associated gravitational-wave signals can provide meaningful constraints on neutron star properties without requiring precise knowledge of the microscopic EoS. At the same time, the systematic deviations observed in exotic matter scenarios offer potential observational signatures that could help discriminate between standard nucleonic models and stars containing dark or strange matter. Future detections of f-mode gravitational waves by next-generation interferometers could therefore serve as a powerful testbed for probing the presence of exotic matter in compact stars and for constraining the strength of dark matter self-interactions. In particular, estimating the signal-to-noise ratios of the predicted f-modes in next-generation detectors such as the Einstein Telescope and Cosmic Explorer, for realistic excitation energies and astrophysical channels (post-merger remnants, proto-neutron stars, or glitches), will be an important direction. Such studies could establish whether the parameter space explored here can be probed observationally, and whether stacking strategies or Bayesian inference using universal relations can enhance detectability. This will be pursued in future work.
Acknowledgments
I sincerely thank Professor Anthony Thomas (University of Adelaide) for his invaluable corrections and insightful suggestions, which have significantly enhanced the clarity and accuracy of this work. His expertise and guidance were instrumental in refining our analysis, and I deeply appreciate his suggestion to this study.

### v3

@@
-Examining dark neutron decay using f-mode gravitational waves
+F-mode Oscillations of Neutron Stars with Dark Matter from Neutron Decay: Implications for Gravitational-Wave Detectability
 Abstract
@@
 It has been shown that neutron stars composed of nucleons-only and self-interacting dark matter exhibit a universal behavior in damping time and angular frequency, whereas strange matter and non-self-interacting dark matter deviate from this trend.
-These findings highlight the significance of gravitational wave detections in investigating neutron star structure, particularly with next-generation observatories such as the Einstein Telescope and Cosmic Explorer.
-In particular, the f-mode frequencies generated by such stars fall within the detection range of modern telescopes, making it possible to observe these signals directly.
-This opens a promising pathway for probing dark matter indirectly—using neutron stars as natural laboratories for fundamental physics.
+Importantly, it has been shown that for a GW energy release of $E\sim 10^{52}$ erg and a source distance of 25 Mpc, the characteristic strain and signal-to-noise ratio exceed the ET-D sensitivity threshold below $\sim$2.1 kHz for all models except the non-interacting DM case, demonstrating that neutron-to-dark matter decay scenarios, including the role of DM self-interactions, can be tested through next-generation gravitational-wave asteroseismology, offering a new probe of DM physics and the neutron lifetime anomaly.
+keywords:
+Neutron stars , Dark Matter , Neutron decay , f-mode oscillations.
+journal: Physics of Dark Universe
+[…]
+organization=School of Science, Technology and Engineering, University of Sunshine Coast, ,addressline=182, Victoria Square, city=Adelaide, postcode=5000, state=SA, country=Australia
 1 Introduction
@@
 This hypothesis implies that neutron stars, composed primarily of neutrons, could serve as natural laboratories for studying dark matter interactions.
-The fundamental mode (f-mode) is a type of non-radial oscillation in compact stars, primarily governed by the star’s bulk properties—such as mass, radius, and internal composition.
+The fundamental mode (f-mode) is a type of non-radial oscillation in compact stars, primarily governed by the star’s bulk properties such as mass, radius, and internal composition.
 These oscillations generate gravitational waves, and their frequency and damping time are sensitive to the equation of state (EoS) of the stellar interior.
@@
 These interdependencies allow the f-mode to act as a bridge between observable gravitational wave signatures and the otherwise inaccessible interior physics of compact stars.
-In this study, the neutron decay into dark matter particle model is tested using well-established universal relations for neutron stars (NSs) involving f-mode properties—relations that have been foundational in the field due to their near-independence from the specific choice of nuclear equation of state (EoS), and which have recently been revisited [28, 29].
-This investigation is further extended to isolated compact stars composed of dark matter (DM), modeled as static, spherically symmetric objects with EoSs for both fermionic and bosonic dark matter particles.
-This paper has been divided into different sections.
-The layout of this study is as follows: Section 2 provides theoretical possibility of neutron decay into dark matter particles, examining potential mechanisms and astrophysical constraints.
-Section 3 provides a comprehensive overview of neutron stars, detailing their formation, key properties, and observed phenomena.
-Section 4 focuses on the crucial concept of the Equation of State (EoS) governing the ultradense matter within neutron stars, exploring theoretical models and their implications.
-Building upon this, Section 5 discusses the fundamental structural equations that describe the equilibrium of neutron stars under their immense gravity.
-Section 6 further elaborates on the Equation of State, potentially exploring specific aspects or advanced considerations.
-Section 7 analyzes the intriguing phenomenon of universalities observed in neutron star properties, highlighting equation-of-state independent relationships.
-This section will be further subdivided, with Section 7.1 specifically focusing on the damping time of neutron star oscillations and its physical mechanisms.
-Section 7.2 will then examine the mass-scaled angular frequency of rotating neutron stars and its significance.
-Section 8 provides a comparative analysis of f-mode frequencies across various neutron star models, offering insights into their structural and dynamical characteristics.
-Section 9 concludes the article by summarizing the key findings and discussing their broader relevance to neutron star physics and gravitational wave astronomy.
+In this work, previous studies of neutron decay into dark matter have been extended by incorporating both nucleonic and strange matter EoS within the quark–meson coupling (QMC) model, and by explicitly evaluating the detectability of f-mode oscillations with next-generation gravitational-wave detectors.
+While earlier works have addressed stability and mass-radius constraints for such stars, a systematic investigation of characteristic strain and signal-to-noise ratios (SNR) in the context of ET-D sensitivity has not yet been carried out.
+These results therefore provide the first quantitative assessment of whether gravitational-wave observations of f-modes can constrain dark matter self-interactions and the presence of exotic matter in neutron stars.
+This paper presents, for the first time, a systematic test of the neutron–to–dark matter decay model using f-mode oscillations and signal-to-noise ratio estimates.
+The study is organized as follows.
+Section 2 introduces the theoretical framework for neutron decay into dark matter and discusses astrophysical constraints.
+Section 3 provides an overview of neutron star formation, structure, and observed phenomena, while Section 4 addresses the equations of state governing ultradense matter.
+Section 5 outlines the stellar structure equations, and Section 6 develops advanced considerations of the EoS.
+Section 7 examines universal relations of f-mode oscillations, with subsections devoted to damping times (7.1) and mass-scaled angular frequencies (7.2).
+Section 8 compares f-mode frequencies predicted by nucleonic, strange matter, and dark matter models.
+Section 9 evaluates detectability by confronting the predicted f-mode signals with the Einstein Telescope sensitivity curve and presenting signal-to-noise ratios.
+Section 10 concludes with a summary of the main results and their implications for neutron star physics, gravitational-wave astronomy, and the search for dark matter signatures.
 2 Neutron decay into dark matter
@@
 At ($\beta$)-equilibrium, the chemical equilibrium equations governing the composition of matter are [40, 110, 74]:
-$\mu_{n}=\mu_{\chi},\quad\mu_{n}=\mu_{p}+\mu_{e},\quad\mu_{\mu}=\mu_{e}$ (4.1)
+$\mu_{n}=\mu_{\chi},\quad\mu_{n}=\mu_{p}+\mu_{e},\quad\mu_{\mu}=\mu_{e}$ (1)
 The energy density of the system, including the dark matter, is given by:
-$\epsilon_{H}=\frac{1}{2}m^{2}_{\sigma}\sigma^{2}+\frac{1}{2}m^{2}_{\omega}\omega^{2}+\frac{1}{2}m^{2}_{\rho}\rho^{2}+\frac{1}{\pi^{2}}\int^{k_{F}^{n}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{p}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk$ (4.2)
-$+\frac{1}{\pi^{2}}\int^{k_{F}^{e}}_{0}k^{2}\sqrt{k^{2}+m_{e}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\mu}}_{0}k^{2}\sqrt{k^{2}+m_{\mu}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\chi}}_{0}k^{2}\sqrt{k^{2}+m_{\chi}^{2}}dk\,.$
-where $m_{i}$ is the mass of corresponding particle (subscripted). The nucleon effective mass, $M_{N}^{*}(\sigma)$, given by [108, 109]
-$M_{N}^{*}(\sigma)=M_{N}-g_{\sigma}\sigma+\frac{d\tilde{(}g_{\sigma}\sigma)^{2}}{2}\,.$ (4.3)
-Here $g_{\sigma}$ is the $\sigma$-nucleon coupling constant in free space calculated in the model. The pressure ($P$) is calculated using
-$P=\int\mu_{f}n_{f}-\epsilon.$ (4.4)
-The nucleons-only EoS is taken from [108] and strange matter EoS is drive based on MIT bag model and the EoS used is given by
-$P=\frac{1}{3}(\epsilon-4B)$ (4.5)
-where $\epsilon$ the energy density of the strange matter and $B$ is the bag constant, taken to be 1014gm/fm3 [113].
+$\epsilon_{H}=\frac{1}{2}m^{2}_{\sigma}\sigma^{2}+\frac{1}{2}m^{2}_{\omega}\omega^{2}+\frac{1}{2}m^{2}_{\rho}\rho^{2}+\frac{1}{\pi^{2}}\int^{k_{F}^{n}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk+$ (2)
+$\frac{1}{\pi^{2}}\int^{k_{F}^{p}}_{0}k^{2}\sqrt{k^{2}+M_{N}^{*}(\sigma)^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{e}}_{0}k^{2}\sqrt{k^{2}+m_{e}^{2}}dk+$
+$\frac{1}{\pi^{2}}\int^{k_{F}^{\mu}}_{0}k^{2}\sqrt{k^{2}+m_{\mu}^{2}}dk+\frac{1}{\pi^{2}}\int^{k_{F}^{\chi}}_{0}k^{2}\sqrt{k^{2}+m_{\chi}^{2}}dk$
+where $m_{i}$ is the mass of the corresponding particle (subscripted), and the details of the last term are given in section  6.1. The nucleon effective mass, $M_{N}^{*}(\sigma)$, given by [106, 107]
+$M_{N}^{*}(\sigma)=M_{N}-g_{\sigma}\sigma+\frac{d\tilde{(}g_{\sigma}\sigma)^{2}}{2}\,.$ (3)
+Here $g_{\sigma}$ is the $\sigma$-nucleon coupling constant in free space calculated in the model and $d$ is the scalar polarizability parameter of the nucleon in the QMC model, which accounts for quadratic corrections to the scalar field coupling [106].
+The pressure ($P$) is calculated using
+$P=\int_{i}\mu_{i}n_{i}-\varepsilon,$ (4)
+where $i=n,p,e,\mu,\chi$. The nucleons-only EoS is taken from [106] and strange matter EoS is drive based on MIT bag model, and the EoS used is given by
+$P=\frac{1}{3}(\varepsilon-4B),$ (5)
+[…]
 5 Structural equations
@@
 The potential presence of dark matter within these stars may influence their structural characteristics, resulting in changes to their equation of state (EoS).
-This study examines methods for integrating dark neutron components into the conventional neutron star EoS and solving the Tolman-Oppenheimer-Volkoff (TOV) [114, 115] equations to assess their impact
+This study examines methods for integrating dark neutron components into the conventional neutron star EoS and solving the Tolman-Oppenheimer-Volkoff (TOV) [112, 113] equations to assess their impact.
 TOV equation
-$\frac{dP}{dr}=-\frac{(\epsilon+P)\frac{d\Phi}{dr}}{r}$ (5.1)
-$\frac{d\Phi}{dr}=\frac{M+4\pi r^{3}P}{r(r-2M)}$ (5.2)
-$\frac{dM}{dr}=4\pi r^{2}\epsilon$ (5.3)
-where $P$ is the pressure, $\epsilon$ is the energy ensity, $\Phi$ is the gravitational potential, $M$ is the mass, and $r$ is the radial coordinate
+$\frac{dP}{dr}=-\frac{(\epsilon+P)}{r}\frac{d\Phi}{dr}$ (6)
+$\frac{d\Phi}{dr}=\frac{M+4\pi r^{3}P}{r(r-2M)}$ (7)
+$\frac{dM}{dr}=4\pi r^{2}\epsilon$ (8)
+where $P$ is the pressure, $\epsilon$ is the energy ensity, $\Phi$ is the gravitational potential, $M$ is the mass, and $r$ is the radial coordinate.
 Perturbation Equations for f-mode Oscillations
-The f-mode oscillations are governed by two coupled differential equations [116, 117, 118]:
-$\frac{dW}{dr}=\frac{d\epsilon}{dP}\left[\omega^{2}r^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W\right]-l(l+1)e^{\lambda}V$ (5.4)
-$\frac{dV}{dr}=2\frac{d\Phi}{dr}V-e^{\lambda}\frac{W}{r^{2}}$ (5.5)
-where $W$ represents the radial displacement function, $V$ is the perturbation of the gravitational potential, $\omega$ is the angular frequency of oscillations, and $l$ is the spherical harmonic index.
+The Lindblom–Detweiler formalism for non-radial oscillations [114] has been adopted in this study.
+Here $\nu(r)$ and $\lambda(r)$ are the static metric functions in the line element $ds^{2}=-e^{\nu}dt^{2}+e^{\lambda}dr^{2}+r^{2}d\Omega^{2}$, and $(X,W,V)$ are perturbation variables defined as in [115].
+This ensures dimensional consistency and reproducibility.
+The f-mode oscillations are governed by two coupled differential equations [116, 117, 118]
+$\frac{dX}{dr}=-\frac{l+1}{r}X+\frac{e^{\lambda/2}}{r}\left[\omega^{2}re^{-\nu/2}V+\frac{d\nu}{dr}W\right],$ (9)
+$\frac{dW}{dr}=-\frac{d\nu}{dr}W-\frac{l(l+1)}{r}V+e^{\lambda/2}X,$ (10)
+where $W(r)$ represents the radial displacement function, $V(r)$ is the metric perturbation function, $\omega$ is the angular oscillation frequency, and $l$ is the spherical harmonic index of the perturbation.
 Boundary Conditions and Numerical Integration
 To solve these equations, boundary conditions are applied:
+  • 1.
     At the center:
-    $W(r)=Ar^{l+1},\quad V(r)=-Ar^{l}/l$ (5.6)
+    $W(r)=Ar^{l+1},\quad V(r)=-Ar^{l}/l$ (11)
+  • 2.
     At the surface:
-    $\omega^{2}R^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W=0$ (5.7)
+    $\omega^{2}R^{2}e^{\lambda}-2\Phi V+\frac{d\Phi}{dr}W=0$ (12)
 The equations are integrated numerically. The eigenvalues of the system correspond to the f-mode frequencies.
 6 Relationship between energy density and pressure
-6.1 Effective Lagrangian for neutron–dark matter conversion
-To model the possibility of neutron decay into dark matter inside neutron stars, we extend the nucleonic sector (described here using the QMC model) by including a dark fermion $\chi$ carrying baryon number.
-In order to reproduce the repulsive self-interaction assumed in the EoS analysis, we introduce a dark vector mediator $V_{\mu}$, analogous to the $\omega$ meson in relativistic mean-field models.
+6.1 Effective Lagrangian for neutron to dark matter decay
+To model the possibility of neutron decay into dark matter inside neutron stars, The nucleonic sector (described here using the QMC model) has been extended by including a dark fermion $\chi$ carrying baryon number.
+In order to reproduce the repulsive self-interaction assumed in the EoS analysis, a dark vector mediator $V_{\mu}$ has been introduced, analogous to the $\omega$ meson in relativistic mean-field models.
 The effective Lagrangian reads
-$\mathcal{L}\;=\;\mathcal{L}_{\rm QMC}[n,p,\sigma,\omega_{\mu},\rho_{\mu}]+\bar{\chi}\,(i\not{\partial}-m_{\chi})\,\chi-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}-g_{\chi}\,\bar{\chi}\gamma^{\mu}\chi V_{\mu}+\varepsilon\,(\bar{n}\chi+\bar{\chi}n),$ (6.1)
+$\mathcal{L}\;=\;\mathcal{L}_{\rm QMC}[n,p,\sigma,\omega_{\mu},\rho_{\mu}]+\bar{\chi}\,(i\not{\partial}-m_{\chi})\,\chi-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}-g_{\chi}\,\bar{\chi}\gamma^{\mu}\chi V_{\mu}+\varepsilon\,(\bar{n}\chi+\bar{\chi}n),$ (13)
 where $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$.
 The last term represents an effective in-medium mixing between neutrons and dark fermions, which allows for neutron conversion into dark matter when $\mu_{n}>\mu_{\chi}$.
-Baryon number is conserved if $\chi$ carries baryon number =1=1.
+Although the effective in-medium operator $\varepsilon\bar{n}\chi+\text{h.c.}$ mixes neutrons with dark fermions, baryon number is conserved if the dark fermion $\chi$ carries baryon number $B(\chi)=1$.
+In this case, the conversion $n\leftrightarrow\chi$ redistributes baryon number between visible and dark sectors but does not violate overall conservation.
+This treatment follows the effective approach proposed in [16, 18, 17].
 At scales below $m_{V}$, integrating out the heavy vector yields a four-fermion operator,
-$\mathcal{L}_{\rm eff}\supset\frac{g_{\chi}^{2}}{2m_{V}^{2}}\,(\bar{\chi}\gamma_{\mu}\chi)(\bar{\chi}\gamma^{\mu}\chi)\equiv\frac{G}{2}\,(\bar{\chi}\gamma_{\mu}\chi)^{2},$ (6.2)
+$\mathcal{L}_{\rm eff}\supset\frac{g_{\chi}^{2}}{2m_{V}^{2}}\,(\bar{\chi}\gamma_{\mu}\chi)(\bar{\chi}\gamma^{\mu}\chi)\equiv\frac{G}{2}\,(\bar{\chi}\gamma_{\mu}\chi)^{2},$ (14)
 where the effective coupling is defined as $G=g_{\chi}^{2}/m_{V}^{2}$.
@@
 The results indicate that the nucleon-only EoS is the stiffest among the considered models.
-As the DM–DM self-repulsion increases, the corresponding EoS for neutron decay into dark matter becomes progressively stiffer and approaches the nucleon-only case which is in agreement with the findings of [50, 41, 51].
+As the DM–DM self-repulsion increases, the corresponding EoS for neutron decay into dark matter becomes progressively stiffer and approaches the nucleon-only case, which is in agreement with the findings of [48, 39, 49].
 This behavior arises because stronger DM–DM self-interaction raises the energy cost for neutron decay into dark matter, thereby reducing the overall dark matter population within the system [48].
@@
 Shown are nucleon-only (QMC), strange matter (MIT bag model), and dark-matter–admixed stars with varying self-interaction strengths $G$.
-The horizontal band indicates the observed $2,M_{\odot}$ pulsar constraint.
-Increasing $G$ stiffens the dark-matter EoS, raising the maximum mass toward the nucleon-only case.
 The radius of light neutron stars is smaller because the EoS is not optimized in the lower energy density region (outer crust region) [107].
@@
 The damping time $\tau$ of f-mode oscillations is related to the balance between the total mode energy $E$ and the rate of energy loss via gravitational-wave emission $\dot{E}{\rm{}_{GW}}$,
-$\tau=\frac{2E}{\dot{E}_{GW}}.$ (7.1)
+$\tau=\frac{2E}{\dot{E}_{GW}}.$ (15)
 The oscillation energy for a mode with angular frequency $\omega$ is given by
-$E=\frac{1}{2}\omega^{2}\int\rho(r)|\xi(r)|^{2}d^{3}r,$ (7.2)
+$E=\frac{1}{2}\omega^{2}\int\rho(r)|\xi(r)|^{2}d^{3}r,$ (16)
 where $\xi(r)$ is the Lagrangian displacement vector and $\rho(r)$ the local energy density. The gravitational-wave luminosity for the dominant quadrupole mode is
-$\dot{E}{\rm{}_{GW}}=-\frac{32G}{5c^{5}}\omega^{6}\left|\int\rho(r)r^{2}Y{2m}(\theta,\phi)drd\Omega\right|^{2}.$ (7.3)
-From these relations, the universal scaling $\tau\sim R^{4}/M^{3}$ emerges [125], which we adopt and extend in this study.
+$\dot{E}{\rm{}_{GW}}=-\frac{32G}{5c^{5}}\omega^{6}\left|\int\rho(r)r^{2}Y{2m}(\theta,\phi)drd\Omega\right|^{2}.$ (17)
+From these relations, the universal scaling $\tau\sim R^{4}/M^{3}$ emerges [125], which has been adopted and extended in this study.
 Reference [126] introduced a first-order universal relation for the damping time ($\tau$) of f-mode oscillations, derived from the quadrupole radiation formula.
 The relation, shown in Figure (3), is expressed as
-$\tau\sim\frac{R^{4}}{M^{3}}.$ (7.4)
+$\tau\sim\frac{R^{4}}{M^{3}}.$ (18)
 Here $\tau$ reflects the ratio of oscillation energy to gravitational-wave luminosity, implying that more massive stars radiate more efficiently and exhibit shorter damping times.
@@
 In comparison, strange matter stars exhibit longer damping times than both nucleon-only stars and neutron stars with dark matter cores.
-Figure 3: The normalised damping time of the f-mode as a function of stellar mass.
-Nucleon-only stars exhibit the shortest damping times, while strange-matter stars show longer-lived oscillations.
-Dark-matter–admixed stars interpolate between these limits depending on $G$.
+Figure 3: The normalised damping time of the f-mode as a function of stellar mass. Nucleon-only stars exhibit the shortest damping times, while strange-matter stars show longer-lived oscillations.
 7.2 Mass-scaled angular frequency
@@
 Figure 5: The relationship between mass-scaled angular frequency and effective compactness is plotted for selected dark matter EoSs, including various DM-DM self-interaction strengths.
-Figures (4) and (5) illustrate the relation between the mass-scaled angular frequency, where $\omega_{r}=2\pi f$, and two structural parameters of neutron stars: compactness ($M/R$) and $\psi=\sqrt{M^{3}/I}$ [117].
+Figures (4) and (5) illustrate the relation between the mass-scaled angular frequency, where $\omega_{r}=2\pi f$, and two structural parameters of neutron stars, namely, compactness ($M/R$) and $\psi=\sqrt{M^{3}/I}$ [117].
 These diagnostics are particularly valuable since they encode information about the distribution of mass and the role of internal composition in determining stellar oscillation properties.
@@
 Figure (5) further reinforces the presence of a quasi-universal relation across different microscopic models when the angular frequency is expressed in terms of $\psi=\sqrt{M^{3}/I}$.
-Remarkably, even when exotic compositions are considered—including strange matter and neutron-decay into DM with strong self-interaction ($G>26$ fm2)—the data points collapse onto a nearly single curve.
+Remarkably, even when exotic compositions are considered, including strange matter and neutron-decay into DM with strong self-interaction ($G>26$ fm2), the data points collapse onto a nearly single curve.
 This indicates that $\psi$ effectively captures the essential structural scaling of neutron stars, smoothing out differences in composition.
@@
 This consistency suggests that such models could be effectively probed through next-generation gravitational wave observatories.
-Our results are consistent with earlier works exploring neutron-to-dark matter decay in neutron stars.
-Motta et al.
-[42] and Grinstein et al.
-[17] found that non-interacting dark fermions destabilize neutron stars, reducing their maximum mass below the $2M_{\odot}$ threshold, which we confirm in our analysis.
-Husain et al.
-[50] similarly emphasized the necessity of DM self-repulsion, consistent with our finding that $G>26$ fm2 restores compatibility with observational constraints.
-For f-mode oscillations, our study extends the universal relations established by Andersson and Kokkotas [125] and refined by Lioutas and Stergioulas [126], showing that these relations persist even in the presence of exotic matter.
-However, we demonstrate that strange matter and non-self-interacting DM produce systematic deviations, suggesting that precise future detections could be used to discriminate between EoSs.
+For f-mode oscillations, this study extends the universal relations established by Andersson and Kokkotas [125] and refined by Lioutas and Stergioulas [127], showing that these relations persist even in the presence of exotic matter.
+However, it has been demonstrated that strange matter and non-self-interacting DM produce systematic deviations, suggesting that precise future detections could be used to discriminate between EoSs.
 Figure 6: The f-mode frequencies are calculated for various equations of state (EoSs), taking into account different dark matter (DM) self-interaction strengths, agreeing to Ref. [128, 127].
+Figure 7 presents the predicted characteristic strain of f-mode oscillations for the considered EoS models, compared with the ET-D sensitivity curve [129, 130].
+The strain values have been calculated assuming a fiducial gravitational-wave energy release of E = 1052 erg, a source distance of D = 25Mpc, and a damping time of $\tau$ = 0.1s, consistent with estimates in [131, 132].
+Under these benchmark conditions, the non–self-interacting dark matter EoS (G0) yields strain amplitudes that remain below the ET sensitivity threshold across the relevant frequency range, implying that such signals would not be detectable.
+By contrast, the self-interacting dark matter and strange matter EoSs predict higher characteristic strains that exceed the ET sensitivity curve for frequencies below 2.1 kHz.
+This demonstrates that, if sources within 25 Mpc deposit of order 1052 erg into f-mode oscillations, the corresponding signals could in principle be detected by ET.
+These results emphasize that f-mode detectability is highly sensitive to the underlying neutron-star EoS, and they suggest that next-generation detectors such as ET can provide meaningful constraints on dark matter interactions through neutron-star asteroseismology.
+Figure 7: Characteristic strain $h_{c}$ of f-mode oscillations from neutron stars as a function of frequency, compared with the sensitivity curve of the Einstein Telescope (ET-D) [130].
+Results are shown for several equations of state (EoS), assuming a total gravitational-wave energy release of E = 1052 erg, damping time = 0.1 sec and a source distance of D = 25 Mpc.
+Figure (8) shows the variation of SNR with f-mode frequency for the different EoS models considered.
+It has been found that the SNR exceeds unity for frequencies below approximately 2.1 kHz.
+This implies that, within this frequency range, f-mode oscillations could be detectable by the Einstein Telescope, provided the assumed energy budget and source distance.
+At higher frequencies, the SNR falls below unity, indicating that detection becomes unlikely due to the reduced strain amplitude and the rising noise floor of the detector.
+Notably, the non–self-interacting DM model (G0) yields systematically lower SNRs, while the interacting dark matter and strange matter EoS predict higher SNR values, suggesting more favorable detectability.
+These results emphasize that the detection prospects of f-mode signals strongly depend on both the EoS and the frequency band of emission, with ET’s sensitivity window favoring signals in the 2.1 kHz regime.
+Figure 8: Signal-to-noise ratio (SNR) of f-mode gravitational-wave signals as a function of frequency, computed for different equations of state (EoS) at a source distance of D = 25 Mpc, damping time = 0.1 sec and energy release E = 1052 erg.
+The detectability of f-mode gravitational-wave signals is highly sensitive to both the neutron star equation of state (EoS) and the frequency band in which the mode energy is emitted.
+The physics behind these trends can be understood as follows.
+Softer equations of state, such as those with exotic matter components, generally produce lower f-mode frequencies, shifting the emission closer to the frequency band where ET is most sensitive.
+Moreover, longer damping times associated with strange matter enhance the overall energy radiated in the mode, further increasing the characteristic strain.
+Conversely, the absence of DM self-repulsion (G0) leads to reduced pressure support, suppressing the excitation energy available to the mode and lowering the strain amplitude.
+Thus, the inclusion of dark matter with repulsive self-interactions not only restores astrophysical stability but also improves the prospects for gravitational-wave detection.
+From an observational standpoint, these results indicate that ET has the potential to place meaningful constraints on the nature of dark matter interactions and the composition of neutron stars.
+A confirmed detection of f-mode oscillations in the $\sim$1–2 kHz range would favor scenarios involving exotic matter or self-interacting dark matter, whereas a systematic absence of such signals could be used to rule out or constrain these models.
+Future work should extend this analysis to population studies, varying source distances, masses, and energy release efficiencies, in order to more comprehensively assess the detectability of f-modes in realistic astrophysical environments.
 9 Conclusion
-In this work, we investigated the impact of nucleonic, strange matter, and dark matter equations of state on the f-mode oscillation properties of neutron stars, with particular attention to damping times and universal relations.
-Our results demonstrate that the nucleon-only EoS is consistently the stiffest, while the inclusion of dark matter generally softens the stellar structure, unless significant DM–DM self-interactions are present.
-Increasing the self-repulsion parameter $G$ suppresses dark matter production, thereby stiffening the EoS and bringing the neutron-decay scenario closer to the nucleonic case.
-In contrast, strange matter exhibits longer damping times and notable deviations at low densities, reflecting its softer pressure support relative to nucleonic matter.
-Importantly, we find that neutron decay into dark matter leads to stellar configurations that remain compatible with current astrophysical constraints.
-Moreover, the corresponding f-mode frequencies of such stars fall squarely within the detection range of modern gravitational-wave interferometers, underscoring their observational relevance.
-Despite these compositional differences, the mass-scaled angular frequency retains quasi-universal behavior when expressed as a function of compactness ($C=M/R$) and the structural parameter $\psi=\sqrt{M^{3}/I}$.
-This universality is preserved even in the presence of exotic matter, with only minor deviations in the case of strange matter and non-self-interacting dark matter.
-The persistence of such relations highlights their robustness and utility as probes of neutron star structure.
-Astrophysically, these results suggest that f-mode oscillations and their associated gravitational-wave signals can provide meaningful constraints on neutron star properties without requiring precise knowledge of the microscopic EoS.
-At the same time, the systematic deviations observed in exotic matter scenarios offer potential observational signatures that could help discriminate between standard nucleonic models and stars containing dark or strange matter.
-Future detections of f-mode gravitational waves by next-generation interferometers could therefore serve as a powerful testbed for probing the presence of exotic matter in compact stars and for constraining the strength of dark matter self-interactions.
-In particular, estimating the signal-to-noise ratios of the predicted f-modes in next-generation detectors such as the Einstein Telescope and Cosmic Explorer, for realistic excitation energies and astrophysical channels (post-merger remnants, proto-neutron stars, or glitches), will be an important direction.
-Such studies could establish whether the parameter space explored here can be probed observationally, and whether stacking strategies or Bayesian inference using universal relations can enhance detectability.
-This will be pursued in future work.
-Acknowledgments
+In this work, the f-mode oscillation properties have been explored of neutron stars containing nucleonic matter, strange matter, and dark matter produced through neutron decay, with particular attention to the role of dark matter self-interactions.
+By computing f-mode frequencies, damping times, characteristic strains, and signal-to-noise ratios, the detectability of these oscillations with the Einstein Telescope (ET) is assessed.
+The nucleonic EoS remains the stiffest, producing systematically higher f-mode frequencies, while the inclusion of dark matter generally softens the stellar structure.
+Strong DM–DM self-repulsion ($G\gtrsim 26$ fm2) counteracts this softening and restores compatibility with astrophysical mass constraints.
+Strange matter predicts longer damping times and lower frequencies, shifting the signal toward the ET’s most sensitive frequency band.
+For an assumed gravitational-wave energy release of $E=10^{52}$ erg, damping time = 0.1 sec and a source distance of $D=25$ Mpc, the characteristic strain and SNR of self-interacting dark matter and strange matter models exceed the ET sensitivity threshold below $\sim 2.1$ kHz.
+In contrast, the non–self-interacting DM case (G0) lies below the sensitivity curve, suggesting that such signals would be undetectable.
+These results indicate that third-generation gravitational-wave observatories could realistically probe exotic matter scenarios through f-mode asteroseismology.
+A confirmed detection of f-mode oscillations in the $\sim$1–2 kHz range would provide strong evidence for exotic matter in neutron star interiors, whereas a lack of such signals could place meaningful constraints on dark matter self-interaction strengths and the viability of strange matter.
+This study demonstrates that f-mode oscillations represent a promising observational probe of neutron star interiors and exotic matter physics.
+The persistence of universal relations across different EoS models highlights their robustness, while the deviations introduced by strange matter and non-self-interacting dark matter suggest possible observational discriminants.
+Future detections of f-mode gravitational waves by instruments such as ET and Cosmic Explorer could therefore open a new window into the microphysics of dense matter, providing unprecedented insight into the role of dark matter and exotic phases in compact stars.
+This analysis extends previous work by explicitly connecting exotic matter EoS models to gravitational-wave observability, making this the first quantitative detectability study of neutron-decay dark matter through f-mode oscillations.
+The results indicate that third-generation interferometers such as ET could provide decisive tests of dark matter self-interactions and the presence of exotic matter in compact stars.
+Acknowledgment
 I sincerely thank Professor Anthony Thomas (University of Adelaide) for his invaluable corrections and insightful suggestions, which have significantly enhanced the clarity and accuracy of this work.